OpenStudy (anonymous):

Yet again find the requested function value meeting all of the given conditions cosO=-√2/2 and tanO<0 find cosO

OpenStudy (anonymous):

find cosO or sinO ?

OpenStudy (anonymous):

My mistake it wants to find cotO

OpenStudy (anonymous):

from the given conditions, the angle in the 2nd quadrant, and by using the phythorian therom ==> 2+n^2=4===> n=sqrt(2) ===> cotO = cosO/sinO = [-sqrt(2)/2]/[sqrt(2)/2] = -1

OpenStudy (anonymous):

so cotO=1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yes it's -1